really need help on this one!
Find the solution to the equation: (5/x+5)+(x/x+3)+(1/x^2+8x+15) = 0
the choice of answers given are: {-4,-3} {-5,1) (-8,-2} and empty set

Answer by stanbon(75887)

Find the solution to the equation: (5/x+5)+(x/x+3)+(1/x^2+8x+15) = 0
the choice of answers given are: {-4,-3} {-5,1) (-8,-2} and empty set
---------------------------
Multiply through be (x+5)(x+3) which is x^2+8x+15, to get:
5(x+3)+x(x+5)+1=0
5x+15+x^2+5x+1=0
x^2+10x+16=0
x=[-10+-sqrt(10^2-4*16)]/2
x=[-10+-sqrt(36)]/2
x=[-10+-6]/2
x=-8 or x=-2
Cheers,
Stan H.